# RNNs, CNNs and Transformers in Human Action Recognition: A Survey and A Hybrid Model

### v1

@@ 3.1 CNN-based approaches in HAR
This section recalls the most prominent CNN-based approaches in HAR, organized chronologically.
@@
Deep learning was still in its early stages in 2012, and CNNs, or CNNs, had not yet gained significant popularity in the field of HAR. The focus was primarily on traditional machine learning approaches, such as Support Vector Machines (SVMs) [49], and handcrafted features, such as Histogram of Oriented Gradients (HOG) [50] and Histogram of Optical Flow (HOF) [51]. A few studies did, nevertheless, start looking into neural networks for action recognition.
In 2012, Ji et al. [52] presented an exploration into the use of 3D CNNs for recognizing human actions in surveillance videos. This model innovatively extends traditional 2D CNNs by performing 3D convolutions, thereby extracting features from both spatial and temporal dimensions to effectively capture motion information. Evaluated on the TREC Video Retrieval Evaluation (TRECVID) dataset and the KTH dataset, the 3D CNN model demonstrated significant improvements over traditional 2D CNNs and other baseline methods in action recognition tasks. The paper’s novelty lies in its unique approach to integrating spatial-temporal features for action recognition, offering enhanced performance in real-world surveillance scenarios. For papers specifically from 2012 that only used CNN architectures for HAR, there may not be a significant number of high-impact papers due to the fact that the use of CNNs in this field was still emerging and not as established as in later years. The seminal works that laid the groundwork for using CNNs in action recognition often came slightly later, as the computational power and dataset availability became more conducive to deep learning approaches.
In 2013, Valle et al. [53] address the challenge of HAR from video data, focusing specifically on the recognition of walking and running actions. They propose an innovative approach using a variant of a 2D CNN, a type of multilayer perceptron, which is distinct for its ability to learn feature extraction directly from raw, unprocessed images during the error backpropagation phase. This methodology represents a significant advancement over traditional methods that rely on pre-processed images. The authors developed a specialized dataset comprising 4000 images per activity to train and evaluate their model. The paper’s contribution lies in its novel approach to feature extraction in HAR, demonstrating the potential of CNNs in efficiently and accurately recognizing complex human actions directly from video sequences, a task of growing importance in fields such as surveillance and automated monitoring systems.
@@
In 2014, the use of CNNs in action recognition was at a pivotal stage, marking a shift from hand-crafted feature-based methods to deep learning approaches. The key points of the use of CNNs in action recognition at that period of time is the following. Emergence of Deep Learning, deep learning, particularly CNNs, had started to dominate image classification tasks, thanks to their ability to learn feature representations directly from raw pixel data. This success in static images paved the way for applying CNNs to video data for action recognition. Challenges in Video data, unlike 2D images, videos incorporate a third dimension which represents the temporal patterns, making action recognition more complex. CNNs had to be adapted to not only recognize spatial patterns but also capture motion information over time dimension. Datasets and Benchmarks, the adoption of large-scale video datasets like UCF-101 [54] and HMDB-51 [55] became more common. These datasets provided diverse sets of actions and were large enough to train deep networks. The performance on these benchmarks has been becoming a key measure of progress for action recognition models. Transfer Learning, due to the computational expense of training CNNs from scratch and the relatively smaller size of video datasets compared to image datasets, transfer learning became a popular strategy. Networks pre-trained on large image datasets like ImageNet [34] were fine-tuned on video frames for action recognition tasks. Computational Constraints, despite the promise of CNNs, computational constraints were a significant challenge. Training deep networks required significant GPU power, and processing video data with CNNs was resource-intensive. This limited the complexity of the models that could be trained and the size of the datasets that could be used.

### v2

@@ 3.1 CNN-Based Approaches in HAR
This section recalls the most prominent CNN-based approaches in HAR based on the model type (i.e., the two-stream CNN, 3D CNN, and RNNs with CNNs), organized chronologically.
@@
Deep learning was still in its early stages in 2012, and CNNs or RNNs had not yet gained significant popularity in the field of HAR. The focus was primarily on traditional machine learning approaches, such as support vector machines [19], and handcrafted features, such as histogram of oriented gradients [21] and histogram of optical flow [7]. A few studies did, nevertheless, start looking into neural networks for action recognition.
@@
In 2014, the use of CNNs in action recognition was at a pivotal stage, marking a shift from hand-crafted feature-based methods to deep learning approaches. The key points of the use of CNNs in action recognition at that period of time are the following. (I) Emergence of deep learning: deep learning, particularly CNNs, had started to dominate image classification tasks, thanks to their ability to learn feature representations directly from raw pixel data. This success in static images paved the way for applying CNNs to video data for action recognition. (II) Challenges in video data: unlike 2D images, videos incorporate a third dimension which represents the temporal patterns, making action recognition more complex. CNNs had to be adapted to not only recognize spatial patterns but also capture motion information over time dimension. (III) Datasets and benchmarks: the adoption of large-scale video datasets like UCF-101 [83] and HMDB-51 [59] became more common. These datasets provided diverse sets of actions and were large enough to train deep networks. The performance on these benchmarks has been becoming a key measure of progress for action recognition models. (IV) Transfer learning: due to the computational expense of training CNNs from scratch and the relatively smaller size of video datasets compared to image datasets, transfer learning became a popular strategy. Networks pre-trained on large image datasets like ImageNet [23] were fine-tuned on video frames for action recognition tasks. (V) Computational constraints: despite the promise of CNNs, computational constraints were a significant challenge. Training deep networks required significant GPU power, and processing video data with CNNs was resource-intensive. This limited the complexity of the models that could be trained and the size of the datasets that could be used.
3.1.1 Two-Stream CNNs